# Batched Self-Consistency Improves LLM Relevance Assessment and Ranking

### v1

@@ 4 Experimental Setup
We evaluate the effects of batching and self-consistency on the relevance assessment and ranking abilities of three LLMs, namely GPT-4o (128K), Sonnet 3 (200K), and Amazon Nova Pro (300K) at temperature 1 across three passage retrieval datasets, releasing anonymized code. [^4] https://anonymous.4open.science/r/batched-sc-emnlp/
@@
The tested number of self-consistency calls per passage, $m$, ranges from 1 (i.e., no self-consistency) to 15. We ensure that for each query, each passage appears in exactly $m$ calls for all LLM methods.
@@ 4.1 Passage Relevance Assessment and Ranking Tasks
We use NDCG@10 to evaluate LLM ranking quality. We treat relevance assessment (PW methods only) as a binary classification task, first converting each LLM score $s_{q,p}\in[0,3]$ to a relevance probability as $p(\hat{y}_{q,p}=1)=s_{q,p}/3$, and then using the area under the precision recall curve (AUC-PR) for evaluation. [^5] AUC-PR is preferred to AUC-ROC for imbalanced data Saito and Rehmsmeier (2015), and IR datasets are highly imbalanced.
@@
To study the effect of the total number of passages evaluated, we test a short ($D=30$) and long ($D=90$) initial list length in what we call shallow and deep search, respectively. All sub-batching methods split the initial list into three equal sized batches (i.e., 10 and 30 passages per batch for shallow and deep search, respectively). The longest resulting batch (90 TREC Covid passages) is roughly 24K words – well within the context limits of all three LLMs.
@@ 4.2 Datasets
Our evaluation includes two well-known open-source datasets, TREC DL-19666https://github.com/microsoft/TREC-2019-Deep-Learning and TREC Covid Thakur et al. (2021) with BM25 used to retrieve the initial passage list $L^{q}$ for each $q$. We also test a third, closed-source dataset from a large legal technology company which we call Legal Search. Legal Search comprises 100 legal queries and passages chunked from 100,000 proprietary legal documents and uses the output of a multi-stage industrial retrieval pipeline to retrieve $L^{q}$. Using both open- and closed-source data allows us to examine whether our findings generalize between data that will have been seen by the LLMs during pretraining and data which has not Törnberg (2024).
@@ 5.1 RQ1: Effects of Batching and Self-Consistency on Relevance Assessment
Figures 2 and 3 and Table 1 show the effects of increasing the number of self-consistency LLM calls/passage ($m$) from 1 (i.e. no self-consistency) to 15 on relevance assessment performance (AUC-PR), with more plots shown in Appendix A. While one-by-one PW methods are competitive at $m=1$, they are always outperformed by a batched PW method at $m=15$.
@@
[…]
Figure 4: Mean relevant and irrelevant LLM scores with sub-batches of 30 psgs/batch (left) vs. all-in-one batches of 90 psgs/batch (right). The sub-batches are relatively consistent in discriminating relevance throughout the batch, while the all-in-one batches lose most of their discriminative power towards the tail of the batch.
@@
At $m=15$, the best performance acoss all datasets and LLMs was always achieved by a batched method with shuffling, which can likely be attributed to more diverse LLM sampling with reduced position bias across self-consistency calls.
Figure 5: Mean LLM scores for one-by-one PW versus batched PW methods with initial order. For all-in-one PW methods, GPT-4o tracks much more closely to one-by-one PW than Sonnet and Nova Pro. Sub-batching with initial order creates artificial cycles and discontinuities at batch junctions.
@@
While all batched self-consistency methods performed well in shallow search (30 passages), for deep search (90 passages), the all-in-one methods performed poorly – far worse than the sub-batched methods (green lines in Figure 3). RQ2 further explores why large batches of 90 passages degraded performance at the values of $m$ tested.
@@ 5.2 RQ2: Effects of Position Biases
Before considering the effect of the initial passage order (RQ3), we first ask whether batched LLM scoring exhibits consistent positional biases (e.g., the first few passages in a batch are scored higher) across random permutations of a given passage list. Addressing this question, Figure 4 shows LLM scores versus passage positions in a batch, with each passage seen by an LLM in $m=15$ random permutations (90 total passages).
@@
The harmful biases in the large 90 passage batches in Figure 4 are obvious: the capacity to discriminate between relevant vs. non-relevant passages is almost gone in the area towards the tail of the batch, with GPT-4o showing a clear lost-in-the-middle effect Liu et al. (2024). By comparison, sub-batching with 30 passages per batch is far more consistent in being able to discriminate relevance throughout the batch, explaining its far superior performance on deep search.
@@ 5.3 RQ3: Effects of Initial Passage Order
Next we investigate the potential positional biases caused by the initial list order $L^{q}$. Figure 5 compares several batched methods that use the initial order to one-by-one LLM scoring, which does not depend on the order of $L^{q}$.
@@
For GPT 4o all-in-one PW in Figure 5, though the front and tail of the batch have slightly higher scores, overall, the batched scores track quite closely to one-by-one PW scores. In contrast, batched Sonnet and Nova Pro scores are typically far lower than their respective one-by-one scores everywhere except at the very front of the batch, helping explain the weakness of these LLMs.
@@
The RHS of Figure 5 shows that sub-batching with initial order can cause score peaks at the start of every sub-batch, creating discontinuities at batch junctions, and inducing cyclical score fluctuations. This explains why sub-batching with the initial order performs worse than the shuffling variants.
@@ 5.4 RQ4: Ranking Performance of Batched Self Consistency Methods
Ranking performance in terms of NDCG@10 for all LLM methods is shown in Tables 2 and 3 for shallow and deep search, respectively, with detailed results on the effects of $m$ in Appendix B.
@@
Without self consistency ($m=1$), one-by-one PW and LW (initial order) methods are competitive rankers, but adding self-consistency helps batched PW methods more than it helps these baselines, making the batched PW methods with $m=15$ the strongest rankers overall. For instance, as seen in Table 3 for Legal Search (deep), GPT-4o one-by-one PW ranking improves from 44.9% NDCG@10 with $m=1$ to 46.8% with $m=15$, while sub-batched (STB) PW ranking improves from 43.8% to 51.3%, respectively (while needing 30 times fewer LLM calls).
@@
When the initial list order $L^{q}$ is kept, LW ranking is competitive with one-by-one PW ranking, but when $L^{q}$ is shuffled, LW methods perform very poorly. Upon examining the LLM outputs, we found that for multiple shuffled input lists $[p_{1},p_{2},...]$ where $p_{i}$ is the $i$’th passage in the batch (i.e., a random passage), the LLM often generated the same rankings (e.g., $[p_{3},p_{12},p_{13},...]$), thus using only the input passage position and not the content. This effect did not occur when initial order was kept.
@@ 6 Conclusion
We propose that batched PW LLM relevance assessment and ranking, where multiple candidates are judged in one LLM call, is especially well-suited for self-consistency. Through multiple experiments, we show that batching amplifies self-consistency benefits, leading to SOTA performance. We conjecture this is because large, diverse input contexts (enabled by candidate batching) increase the diversity of LLM sampling over multiple self-consistency calls.
@@
The main limitation of our work are the computational resources required, since LLM relevance assessment and ranking is expensive computationally. We thus only tested a range of $m\in\{1,\cdots,15\}$ for the number of self-consistency calls, even though higher levels of $m$ would have added information to our results. Also due to computational limitations, we only tested three LLMs (GPT-4o, Claude Sonnet 3, and Amazon Nova Pro) across three datasets (TREC DL-19, TREC Covid Thakur et al. (2021), and Legal Search), though it would be interesting to test an even wider range of models and datasets. Simlarly, we were limited to testing four batch sizes $\in\{1,10,30,90\}$ across five batching strategies (c.f. Figure 1) and two levels of search: shallow (30 initial passages) and deep (90 initial passages).
@@
Appendix A Appendix A: Relevance Assessment Quality vs Scores per Passage
Figures 6 and 7 bellow show the effects of $m$ on AUC-PR of one-by-one PW and batched PW methods for all datasets and LLMs for shallow and deep search, respectively.
@@ Appendix A Appendix A: Relevance Assessment Quality vs Scores per Passage
Appendix B Appendix B: NDCG@10 vs Scores per Passage
Figures 8 and 9 bellow show the effects of $m$ on NDCG@10 of one-by-one PW and batched PW methods for all datasets and LLMs for shallow and deep search, respectively.
@@ Appendix B Appendix B: NDCG@10 vs Scores per Passage
Appendix C Appendix C: Prompt Templates
Figures 10 and 11 show the full prompts used for our PW and LW implementations.

### v2

@@ 4 Experimental Setup
We evaluate the effects of batching and self-consistency on the relevance assessment and ranking abilities of three LLMs, namely GPT-4o (128K), Sonnet 3 (200K), and Amazon Nova Pro (300K) at temperature 1 across three passage retrieval datasets, releasing code. [^4] https://anonymous.4open.science/r/batched-sc-emnlp/
@@
We vary the number of self-consistency calls per passage, $m$, from 1 (i.e., no self-consistency) to 15. For each query, each passage appears in exactly $m$ calls.
@@ 4.1 Passage Relevance Assessment and Ranking Tasks
We use NDCG@10 to evaluate LLM ranking quality. We treat relevance assessment as a binary classification task, converting each LLM score $s_{q,p}\in[0,3]$ to a relevance probability $p(\hat{y}_{q,p}=1)=s_{q,p}/3$, and using the area under the precision recall curve (AUC-PR) for evaluation. [^5] AUC-PR is preferred to AUC-ROC for imbalanced data Saito and Rehmsmeier (2015), and IR datasets are highly imbalanced.
@@
To study the effect of the total number of passages evaluated, we test a short ($D=30$) and long ($D=90$) initial list length in what we call shallow and deep search, respectively. All sub-batching methods split the initial list into three equal sized batches (i.e., 10 and 30 passages per batch for shallow and deep search, respectively).
@@ 4.2 Datasets
Our evaluation includes two well-known open-source passage retrieval datasets, TREC DL-19 Craswell et al. (2019) and TREC Covid Voorhees et al. (2021) with BM25 used to retrieve the initial passage list $L^{q}$ for each $q$. We also test a third, closed-source dataset from Thomson Reuters which we call Legal Search. Legal Search comprises 100 legal queries and paragraph-size passages chunked from 100,000 proprietary legal documents and uses the output of a multi-stage industrial retrieval pipeline to retrieve $L^{q}$.
Token Usage
Table 1 shows the mean input token counts per LLM call for each method. The longest resulting input context (90 TREC Covid passages) is 30K-34K tokens, which is well within context limits for all three LLMs.
@@ 5 Experimental Results
Legal Search DL-19 Covid
GPT-4o Sonnet Nova GPT-4o Sonnet Nova GPT-4o Sonnet Nova
Psg/ Psg AUC1/ AUC1/ AUC1/ AUC1/ AUC1/ AUC1/ AUC1/ AUC1/ AUC1/
Batch Order AUC15 AUC15 AUC15 AUC15 AUC15 AUC15 AUC15 AUC15 AUC15
Shallow (30 psg)
1 32 / 41 24 / 28 24 / 29 39 / 52 27 / 31 29 / 33 70 / 78 68 / 73 68 / 77
10 Init. 31 / 44 22 / 27 27 / 33 34 / 51 25 / 33 32 / 40 70 / 77 63 / 70 68 / 74
(sub-batch) STB 31 / 46 22 / 38 25 / 39 39 / 57 25 / 41 30 / 47 69 / 80 65 / 77 68 / 79
BTS 31 / 45 23 / 35 25 / 36 39 / 55 25 / 41 32 / 46 69 / 79 63 / 75 65 / 78
30 Init. 32 / 45 23 / 27 27 / 37 46 / 55 26 / 35 30 / 44 69 / 78 63 / 69 63 / 74
(all psgs) Shuf. 30 / 45 20 / 32 24 / 38 37 / 58 24 / 41 31 / 53 69 / 80 58 / 74 65 / 77
Deep (90 psg)
1 20 / 29 15 / 20 15 / 20 34 / 51 23 / 28 25 / 31 63 / 72 62 / 68 62 / 70
30 Init. 20 / 34 13 / 21 18 / 28 30 / 49 20 / 30 29 / 40 64 / 75 55 / 67 60 / 72
(sub-batch) STB 20 / 36 14 / 27 17 / 31 32 / 52 19 / 40 23 / 42 64 / 77 56 / 73 60 / 75
BTS 22 / 36 14 / 26 17 / 29 29 / 52 18 / 37 25 / 44 64 / 77 55 / 72 60 / 75
90 Init. 16 / 26 11 / 16 13 / 20 22 / 48 16 / 33 23 / 37 52 / 67 48 / 57 45 / 54
(all psg) Shuf. 14 / 29 11 / 21 13 / 20 22 / 50 15 / 42 19 / 44 53 / 73 45 / 65 46 / 60
Table 2: AUCm, representing the AUC-PR at $m$ self-consistency calls/passage, for PW relevance assessment methods at $m=1$ (no self-consistency) and $m=15$. Increasing $m$ improves all PW methods, but the batched PW methods improve faster, becoming the best methods at $m=15$, likely because they create more diverse self-consistency ensembles. The highest AUC-PR for each $m$ and LLM is in bold.
@@ 5.1 RQ1: Effects of Batching and Self-Consistency on Relevance Assessment
Figures 2 and 3 and Table 2 show the effects of increasing the number of self-consistency LLM calls/passage ($m$) from 1 (i.e., no self-consistency) to 15 on relevance assessment performance (AUC-PR) – with more plots in Appendix A. While one-by-one PW methods are competitive at $m=1$, they are always outperformed by a batched PW method at $m=15$.
@@
Though increasing the number of self-consistency calls greatly improved all PW methods, batched methods improved considerably more than one-by-one methods. We conjecture that this is because batching leads to more diverse LLM sampling across self-consistency calls. For instance, for GPT-4o on Legal Search (Shallow), one-by-one PW improved from 32% AUC-PR at $m=1$ to 41% at $m=15$ ($9\%$), while all-in-one PW (Shuffled) improved from 30% to 45% ($15\%)$, respectively.
Batching can reduce latency by an order of magnitude or more:
Figure 4 shows per-query runtimes across different values of $m$ and varying numbers of parallel LLMs for shallow Legal Search. The median time for 30 sequential one-by-one PW LLM calls was 34–58s, depending on the LLM, while judging all 30 passages in a single call took only 2–6s, constituting a 6-17x speedup. Such speedups can be expected if the number of parallel LLMs is not greater than $m$ – otherwise, smaller batch sizes allow for more parallelization, as shown in the second pane of Figure 4. But if there are only a few parallel LLMs available, batched methods are much faster, especially when self-consistency ensembling is used.
@@
At $m=15$, the best performance across all datasets and LLMs was always achieved by a batched method with shuffling, which can likely be attributed to more diverse LLM sampling across self-consistency calls.
@@
While all batched self-consistency methods performed well in shallow search (30 passages), for deep search (90 passages), the all-in-one methods performed poorly – far worse than the sub-batched methods (green lines in Figure 3). RQ2 further explores why the overly large batches of 90 passages degraded performance at the values of $m$ tested.
Figure 5: Mean relevant and irrelevant LLM scores with sub-batches of 30 psgs/batch (left) vs. all-in-one batches of 90 psgs/batch (right). The 30 passage batches are relatively consistent in discriminating relevance throughout the batch, while the 90 passage batches lose most of their discriminative power towards the tail of the batch.
@@ 5.2 RQ2: Effects of Position Biases
Before considering the effect of the initial passage order (RQ3), we first ask whether batched LLM scoring exhibits consistent positional biases across random permutations of a given passage list. Figure 5 shows LLM scores versus passage positions in a batch, with each passage seen by an LLM in $m=15$ random permutations.
@@
The harmful biases in the large 90 passage batches in Figure 5 are obvious: the capacity to discriminate between relevant vs. non-relevant passages is almost gone towards the tail of the batch, with GPT-4o showing a clear lost-in-the-middle effect Liu et al. (2024). By comparison, sub-batching with 30 passages per batch is far more consistent in being able to discriminate relevance throughout the batch, explaining its far superior performance on deep search.
@@ 5.3 RQ3: Effects of Initial Passage Order
Next we investigate the potential positional biases caused by the initial list order $L^{q}$. Figure 6 compares several batched methods that use the order of $L^{q}$ versus one-by-one scoring, which does not depend on the order of $L^{q}$.
Figure 6: Mean LLM scores for one-by-one PW versus batched PW methods with initial order. For all-in-one PW methods (left), GPT-4o tracks much more closely to one-by-one PW than Sonnet and Nova Pro. Sub-batching with initial order (right) creates artificial cycles and discontinuities at batch junctions.
@@
For GPT 4o all-in-one PW in Figure 6, the batched scores track quite closely to one-by-one PW scores, though the front and tail of the batch have slightly higher scores. In contrast, batched Sonnet and Nova Pro scores are typically far lower than their respective one-by-one scores everywhere except at the very front of the batch, helping explain the weakness of these LLMs.
@@
The RHS of Figure 6 shows that sub-batching with initial order can cause score peaks at the start of every sub-batch – creating discontinuities at batch junctions and inducing cyclical score fluctuations. This explains why sub-batching with the initial order performs worse than the shuffling variants.
@@ 5.4 RQ4: Ranking Performance of Batched Self Consistency Methods
Ranking performance in terms of NDCG@10 for all LLM methods is shown in Tables 3 and 4 for shallow and deep search, respectively, with detailed results on the effects of $m$ in Appendix B.
@@
Without self consistency ($m=1$), one-by-one PW and LW (initial order) methods are competitive rankers, but adding self-consistency helps batched PW methods more than it helps these baselines – making the batched PW methods with $m=15$ the strongest rankers overall. For instance, as seen in Table 4 for Legal Search (deep), GPT-4o one-by-one PW ranking improves from 44.9% NDCG@10 with $m=1$ to 46.8% with $m=15$, while sub-batched (STB) PW ranking improves from 43.8% to 51.3%, respectively.
@@
When the initial list order $L^{q}$ is kept, LW ranking is competitive with one-by-one PW ranking, but when $L^{q}$ is shuffled, LW methods perform very poorly.
@@ 6 Conclusion
We show that batched PW methods for passage relevance assessment and ranking can improve not only efficiency, but also judgment quality by enabling content from multiple passages to be seen jointly. Further, when self-consistency ensembles are used to collect and aggregate multiple scores per passage, batched methods can create more diversity between ensemble components than one-by-one methods. While score variation in one-by-one methods comes only from LLM stochasticity, batching can naturally diversify the contexts in which a passage is scored through different co-candidate subsets and permutations – with autoregressive multi-score sequence generation diversifying scores even further. As our experiments show, this leads to batching amplifying the test-time scaling benefits of self-consistency, giving batched PW methods the best performance while achieving order-of-magnitude level speedups over one-by-one PW methods.
@@
The main limitation of our work are the computational resources required, since LLM relevance assessment and ranking is expensive computationally. We thus only tested a range of $m\in\{1,\cdots,15\}$ for the number of self-consistency calls, even though higher levels of $m$ would have added information to our results. Also due to computational limitations, we only tested three LLMs (GPT-4o, Claude Sonnet 3, and Amazon Nova Pro) across three datasets (TREC DL-19, TREC Covid, and Legal Search), though it would be interesting to test an even wider range of models and datasets. Similarly, we were limited to testing four batch sizes $\in\{1,10,30,90\}$ across five batching strategies (c.f. Figure 1) and two levels of search: shallow (30 initial passages) and deep (90 initial passages).
@@
Appendix A Relevance Assessment Quality vs Scores per Passage
Figures 7 and 8 below show the effects of $m$ on AUC-PR of one-by-one PW and batched PW methods for all datasets and LLMs for shallow and deep search, respectively.
@@ Appendix A Relevance Assessment Quality vs Scores per Passage
Appendix B NDCG@10 vs Scores per Passage
Figures 9 and 10 below show the effects of $m$ on NDCG@10 of one-by-one PW and batched PW methods for all datasets and LLMs for shallow and deep search, respectively.
Appendix C Experiment Runtimes
Figures 11-15 show the per-query runtimes for all LLM methods for several values of $m$ and numbers of parallel LLMs available.
Appendix D Prompt Templates
Figures 16 and 17 show the full prompts used for our PW and LW implementations.
@@ Appendix D Prompt Templates
Figure 11: Median per-query runtimes for Covid, Shallow (30 total passages), with error bars showing the IQR.
Figure 12: Median per-query runtimes for DL-19 Search, Shallow (30 total passages), with error bars showing the IQR.
Figure 13: Median per-query runtimes for Legal Search, Deep (90 total passages), with error bars showing the IQR.
Figure 14: Median per-query runtimes for Covid, Deep (90 total passages), with error bars showing the IQR.
Figure 15: Median per-query runtimes for DL-19, Deep (90 total passages), with error bars showing the IQR.